OpenStudy (anonymous):

please helppppppp. What is the equation of y = x3 with the given transformations? horizontal translation right 10 units and vertical translation down 12 units A. y = (x - 10)3 - 12 B. y = (x + 10)3 - 12 C. y = (x - 12)3 + 10 D. y = (x + 12)3 + 10

OpenStudy (anonymous):

@bibby

OpenStudy (triciaal):

horizontal changing x left add and right subtract from "base" at 0 vertical adjusts the y value positive add up negative subtract down

OpenStudy (triciaal):

big hint 12 down = -12 so start with A or B

OpenStudy (anonymous):

im sorry, im really confused! any other helpful tips. i learn best if the steps are broken down.

OpenStudy (bibby):

if you add to the base it move the function to the left (i.e. (x+2)^2 moves 2 units left) and moving the base to the right with subtraction

OpenStudy (anonymous):

do i need to find the y & x intercept ? or am i totally off?

OpenStudy (bibby):

that's kinda unnecessary, these questions are more about function transformations. i.e. how do I push this funciton 10 units up or 3 units right or something gimme a few mins to clean up this mess

OpenStudy (anonymous):

okay so i made my x & y chart, and it falls to left and rises right, and i graphed it.

OpenStudy (xapproachesinfinity):

10 units right mean y=(x-10)^3 and 12 units down y=(x-10)^3-12

OpenStudy (anonymous):

okay, where getting somewhere. but how am i suppose to go about solving this.. im still confused.

OpenStudy (bibby):

moving up and down you modify the whole function f(x) (i.e. f(x)+10, f(x)-10) moving right and left means you modify the base f(x+2) or f(x-5))

OpenStudy (xapproachesinfinity):

well it is already solved what else you need?

OpenStudy (anonymous):

how you solved it, i understand u gave me the answer.. but how did you get it?

jimthompson5910 (jim_thompson5910):

Cubic functions are in this general form